Jacob is a member of WCC (an LLC taxed as a partnership). Jacob was allocated $120,000 of business income from WCC for the year. Jacob’s marginal income tax rate is 37 percent. The business allocation is subject to 2.9 percent of self-employment tax and 0.9 percent additional Medicare tax. (Round your intermediate calculations to the nearest whole dollar amount.)

What is the amount of tax Jacob will owe on the income allocation if the income is qualified business income (QBI) and Jacob qualifies for the full QBI deduction?

(I keep getting $39,136 but it's telling me it's the wrong answer)
Answer #1

Business Income allocation:

Note: Self employed taxpayers can deduct employer's contribution on ta, which is 7.65

= 100 - 7.65 = 92.35

Deduction of 50 percent of SE tax = (1) * 0.9235 * 0.029 * 0.5 = $1339

Business Income Allocation $120000

QBI deduction 0 (Income does not qualify for QBI deduction)

Deduction for 50 percent of SE tax (note 1) $(1339)

Net income taxable business income    $1188661

Income tax paid by owner $439804.57 (1188661 * 0.37)

Self employment tax $3213.78 (120000*0.9235*0.029)

Additional medicare tax $997.38 ( 120000* 0.9235 * 0.09)

Owner tax $452992.15

B. Jacob qualifies for QBI deduction:

Business Income Allocation $120000

QBI deduction $(24000) (120000 * 20%)

Deduction for 50 percent of SE tax (note 1) $(1339)

Net income taxable business income    $94661

Income tax paid by owner $35024.57 (94661 * 0.37)

Self employment tax $3213.78 (120000*0.9235*0.029)

Additional medicare tax $997.38 ( 120000* 0.9235 * 0.09)

Owner tax $39235.73